# What is emergence, after all?

### v2

@@ Abstract
The term emergence is increasingly used across scientific disciplines to describe phenomena that arise from interactions among a system’s components but cannot be readily inferred by examining those components in isolation. While often invoked to explain higher-level behaviors—such as flocking, synchronization, or collective intelligence—the term is frequently used without precision, sometimes giving rise to ambiguity or even mystique. In this perspective paper, we clarify the scientific meaning of emergence as a measurable, physically grounded phenomenon. Through concrete examples—such as temperature, magnetism, and herd immunity in social networks—we review how collective behavior can arise from local interactions that are constrained by global boundaries. By disentangling emergence from vague overuse, we emphasize its role as a rigorous tool for understanding complex systems. Our goal is to show that emergence, when properly framed, offers not mysticism but insight.
@@
This manuscript was compiled on August 11, 2025
@@
Emergence occurs where a whole exhibits properties absent from its individual parts—an idea that traces back to Aristotle’s dictum that “the whole is something besides the parts” (1). In science, the term refers to patterns, structures, or behaviors that are not straightforwardly deducible from components in isolation (2, 3, 4, 5, 6, 7), yet are often captured by coarse, predictive descriptions that do not track every micro-detail. You can see emergence in the wild—flocks wheeling as one, fireflies locking into rhythm, insect societies coordinating without commanders (8)—and on through computer simulations, where minimal rules produce striking order: Reynolds’ “boids” (9), cellular automata (10), and the Game-of-Life glider (11).
Philosophical roots of the concept trace back to Mill’s distinction between additive “mechanical” effects and qualitatively new “chemical” ones (12). Lewes later coined the term “emergent” for such cases (13), and early British emergentists (14, 15, 16) developed it into a layered view of nature in which higher-level laws complement, and sometimes influence, lower-level processes. This tradition faded as reductionism gained ground, driven by the intertheoretic reduction models (17, 18, 19). The debate revived with arguments for multiple realizability (20, 21, 22) and Anderson’s claim that “more is different” (23), followed by work on supervenience and causal autonomy (24). In recent years, the term has evolved from describing specific transitions, such as the “emergence of metabolism” (25), to becoming a popular buzzword, gaining more usage in scientific and non-scientific discourses (26, 23). Fig. 1 shows this trend.
Here we re-anchor emergence in measurable, physically grounded mechanisms. We begin with textbook cases from physics, where coarse-graining and symmetry breaking compress microscopic dynamics into macroscopic laws. We then turn to biology and the social sciences, where collective behavior arises from local interactions. Finally, we consider emergence in networks, showing how the same conceptual toolkit applies across domains. We aim to show that emergence is about an economy of description that discards inessential detail while preserving predictive power.
@@
Level of Description
Knowledge unfolds across successive levels of abstraction—from raw experience to language and higher-order concepts (28). Confusion between these levels distorts understanding. Emergence is everywhere, and the surprise it evokes often reflects where we choose to stop asking more questions (29, 24). What we call “emergent” typically depends on the limits of our knowledge, tools, or perspective—in other words, on epistemology (30, 31). Mixing red and green paint yields a dull brown—basic chemistry explains it well enough for most, and few would call the result emergent. Yet a physicist who tracks the same blend down to quantum-mechanical interactions would hardly find the task trivial.
Every phenomenon can be analyzed at multiple descriptive levels, and what feels emergent at one level may be obvious at another (30). Explanation is always a matter of perspective: an economist analyzing a financial crisis need not invoke atoms or molecules but instead works at the social scale relevant to the questions at hand. Even so, emergent behaviors are not merely in the eye of the beholder—they can display objective, quantifiable signatures such as information flow or causal strength (32, 33). We can make emergence more rigorous by formalizing it through coarse-graining maps that discard detail yet preserve prediction.
Figure 2: Coarse-graining with different maps.
(a) A global coarse-graining that yields temperature $T$ by averaging the kinetic energy of all particles, assuming local equilibrium; $\mathcal{F}:\bigl{(}v_{1},\cdots,v_{N}\bigr{)}\;\longmapsto\;T.$ (b-c) Local majority-vote map applied to blocks of four spins $(s_{i},s_{j},s_{k},s_{l})$, returning a block spin $S_{I}$ defined as: $S_{I}=\operatorname{sign}\left(s_{i}+s_{j}+s_{k}+s_{l}-2\right)$, where $\operatorname{sign}(x)$ is 1 for $x>0$, and -1 otherwise; $\mathcal{F}:\bigl{\{}(s_{i},s_{j},s_{k},s_{l}\}\bigr{)}\;\longmapsto\;\{S_{I}\}.$ (c-d) The same map is applied to the previously coarse-grained block. Applying $\mathcal{F}$ on block (d) will result in a single upward block spin.
@@
Emergence is present when there exists a many-to-one map from a micro-level theory (more fundamental, detailed, or lower-level) to a macro-level theory (higher-level), such that the macro description remains predictive even after discarding most of the microscopic detail (34). As Fig. 2 presents, a map is a selective, structured, purpose-driven abstraction—useful precisely because it is not the whole territory it depicts (28, 35). Every map omits, distorts, or abstracts; acknowledging those sacrifices is part of scientific hygiene (35). Coarse-graining leaves us with just enough structure and real patterns to construct an autonomous and useful theory (36). This is why engineers who build bridges do not need to take a course in quantum field theory, or why airplanes can fly safely despite our incomplete understanding of quantum gravity.
Coarse-graining is the process of identifying which features of a system are essential for capturing its macroscopic behavior and which can be systematically discarded. Let the system’s state at time $t$ be given by the vector $\mathbf{x}(t)=\bigl{(}x_{1}(t),\dots,x_{N}(t)\bigr{)}$, where a high-resolution dataset comprises $N$ time series, and each $x_{i}(t)$ represents the evolution of the $i$-th microscopic variable. We can then formally view coarse-graining as a process of lossy data compression, which can be described as a mapping:
@@
where $N^{\prime}<N$, and the coarse-grained observables $\{X_{j}\}$ may be defined on a different (typically coarser) time grid $\{t^{\prime}_{\ell}\}$. Fig. 2 shows two different sets of coarse-graining where they simply average out the observables. The key practical questions here are whether there exists a finite set of macroscopic parameters that can adequately describe the phenomena of interest within a specified tolerance, and if so, which mapping $\mathcal{F}^{\star}$ minimizes the chosen distortion measure while achieving that macro-level description. The outcome is a systematic and metric-aware approach to transition from micro-level dynamics to macro-level dynamics.
The very notion of a map between levels of description tempts us to believe that deeper, more fundamental layers offer better explanations—an intuition reinforced by Physics’ success in reducing systems to atoms and beyond. But how far should we go? What counts as the most fundamental level?
@@
Physics is expressed in languages that are intrinsically redundant—coordinates, gauge potentials, duality frames, and so on—and those redundancies shift as we move between scales (37). What we call “fundamental”—even spacetime and gauge fields—may not exist in a more microscopic theory, but rather emerge after coarse-graining. Spacetime geometry and general relativity might also be emergent, arising from quantum theory through entanglement and boundary dynamics (37). This realization, however, does not make emergence a “secret sauce.” It simply shows how nature organizes complexity through scale, and “fundamental” is often a provisional concept.
@@ A New Ontology
Classical Mechanics itself emerges from a more fundamental quantum theory, which is, in turn, ontologically different. In Classical Mechanics, a particle is a point in phase space with well-defined position and momentum. In Quantum Mechanics, the ontology shifts: a particle is represented by a vector in Hilbert space, and is described by a wavefunction (44). Position and momentum are no longer coordinates but observables. The uncertainty principle ensures that no wavefunction can define both position and momentum with perfect precision. Thus, quantum mechanics offers no definitive answer to the classical question of “where” a particle is, or “how fast” it is moving (44). Nevertheless, the classic macroscopic theory remains applicable in its domain, and it is consistent with the underlying quantum description through the processes of decoherence and measurement, which effectively produce the appearance of wavefunction collapse (45).
@@
Since the advent of Statistical Mechanics, we have learned how coarse-graining microscopic degrees of freedom gives rise to macroscopic quantities like temperature or pressure, and how these quantities are connected through equations of state (46). Statistical mechanics not only explains why thermodynamics works but also defines its domain of applicability. For example, the domain of applicability of temperature can be expressed as
@@
where $m$ is the particle mass, $\langle v^{2}\rangle$ is the ensemble-averaged squared speed, $N$ is the number of particles, and the Boltzmann constant is set to one, $k_{B}=1$ (47). For large $N$, the fluctuation term $1/\sqrt{N}$ vanishes, making temperature a sharply defined coarse-grained variable. For small systems, however, fluctuations are significant, and temperature loses descriptive power. Therefore, in the large-size limit, Thermodynamics is not merely an approximation to statistical mechanics—it is an effective theory (48), one that captures the essential behavior of macroscopic systems using a few key variables at the right scale (49, 50). Instead of tracking the detailed positions and velocities of every particle, we use variables like temperature, pressure, and entropy, governed by their own simple and powerful laws at the level of macrostates (51, 52). Effective theories isolate what matters at a given scale and discard what does not (53). They are not shortcuts; they are self-contained, predictive, and often universal descriptions of emergent levels of reality. A mean-field theory is the zeroth-order effective description (54). It is invaluable as a first cut, exact when fluctuations vanish, but insufficient whenever correlations hold the key to collective behavior (55).
@@
Yet not all effective theories are created equal. In some systems—particularly those involving biological, cognitive, or social dynamics—the coarse-grained variables do not merely summarize the current state of the underlying components. They also retain memory, storing information about past configurations. This historical dependence alters the structure of the effective theory, introducing higher-order terms that reflect feedback, path dependence, and self-reference (34). Such systems are still governed by physical laws, but their dynamics can become computationally intractable, meaning that no tractable microscopic derivation will fully recover the macroscopic behavior (56). The emergence of life, mind, or social structure appears puzzling not because they break with physicalism, but because they represent a qualitatively more complex class of emergent phenomena (56). The first step toward understanding how emergence occurs is to identify when and under what conditions it arises, as well as its immediate consequences.
Onset of Emergence & Symmetry Breaking
@@
A clear and intuitive example of how emergence happens is magnetization. In a fridge magnet, the collective alignment of billions of electron spins produces a macroscopic magnetic property, denoted $m(T)$, which appears only when the system is below a critical temperature $T_{c}$. For a fridge magnet, it is around $450^{\circ}$C. Heating the material beyond this point destroys the alignment, and the magnetic property disappears in a continuous phase transition (47). A single governing equation (Hamiltonian) describes both magnetized and non-magnetized states, and at the critical point $T=T_{c}$, the spin-flip symmetry of the Hamiltonian remains intact, but the actual state chooses a specific direction, resulting in $m\neq 0$ (54). This symmetry breaking defines an order parameter that compresses the full $10^{23}$-spin microstate into a single coarse-grained vector whose dynamics—like domain walls or spin waves—obey new effective laws (57, 58, 23).
@@
At the critical point, key properties of the system—such as correlation length and relaxation time—diverge (59). These divergences, characterized by critical exponents, serve as clear markers of the onset of emergence. More interestingly, different systems composed of distinct elements can exhibit the same critical behavior (60). That is, systems with very different microstructures can fall into the same universality class and be described by the same macroscopic theory. This substrate-independence also means that from the perspective of the emergent behavior, one cannot tell which microscopic system produced it. The liquid–gas transition belongs to the same universality class as the ferromagnetic transition. Universality classes, therefore, organize not just critical exponents but also the very symmetry content that survives at macroscopic scales, reinforcing why broken symmetry is the natural language of emergence (23, 37). In this sense, “more is different” (23): once many degrees of freedom lock together, they generate effective laws—and sometimes entirely new effective symmetries—that are absent from, yet fully compatible with, the underlying dynamics.
Broken symmetries underpin certain emergent behaviors (23, 61, 62, 63), but not all. For example, it is not straightforward to extend the statistical mechanics framework of critical phenomena to describe abrupt qualitative changes in dynamical systems. Likewise, the emergence of life from chemical interactions does not neatly correspond to a phase transition in the physical sense, despite claims to the contrary by some physicists (64). Intricate phenomena such as consciousness or thought could, in principle, arise from entirely different physical substrates—biological neurons or silicon circuits alike (65, 66)—provided there is sufficient evidence that they result from well-defined critical phenomena (67).
Universality & Dualities
@@
Technically, critical universality manifests only in the critical region around a continuous phase transition: as the system approaches the transition, the renormalization-group flow is drawn toward a scale-invariant fixed point that is attractive along irrelevant directions but unstable along relevant ones, making macroscopic behavior largely independent of microscopic details (61). However, this is not the only notion of equivalence and universality in Physics.
Dualities show that theories with ontologically distinct variables—bosons versus fermions, or gravity versus no gravity—can represent the same underlying quantum physics. In $1+1$ dimensions, the bosonic sine–Gordon model is exactly dual to the fermionic Thirring model, exchanging solitonic waves for interacting particles (68). In AdS/CFT, a conformal field theory without gravity on a $d$-dimensional boundary is holographically equivalent to a $(d+1)$-dimensional bulk theory with gravity (69). Even composite objects such as hydrogen atoms—each built from two spin-$\frac{1}{2}$ fermions—can act as effective bosons. Therefore, the large-scale description is not dictated by microscopic “stuff,” but by deeper symmetries and consistency conditions, reminding us that macro structures can arise as alternative micro read-outs of the same physical substrate.
@@
Emergence in Networks
@@
To move beyond examples from physics, we now examine how large-scale patterns can emerge from local interactions in social and informational systems. Consider a set of $N$ points where each pair is connected with probability $p$. This defines an Erdős–Rényi random graph with $N$ nodes and roughly $p\binom{N}{2}$ links or connections. As $p$ increases, the network shifts from isolated nodes ($p=0$) to a fully connected graph ($p=1$), with clusters forming in between. Given the value of $p$, it is interesting to look at the behavior of the cluster with the most number of connected nodes, the giant component. Letting $\sigma$ be the fraction of nodes it contains, we find that as in the large-size limit ($N\to\infty$), the network undergoes a phase transition such that the size of its giant component is zero below a critical value $p_{c}$, and positive above it. Here, $p$ is the control parameter, and $\sigma$ represents the order parameter, mirroring magnetization in ferromagnets, which appears below the critical temperature. When $p>p_{c}$, large-scale connectivity emerges, and the network is said to percolate (70).

### v3

@@ Abstract
The term emergence is increasingly used across scientific disciplines to describe phenomena that arise from interactions among a system’s components but cannot be readily inferred by examining those components in isolation. While often invoked to explain higher-level behaviors—such as flocking, synchronization, or collective intelligence—the term is frequently used without precision, sometimes giving rise to ambiguity or even mystique. In this perspective paper, we clarify the scientific meaning of emergence as a measurable and physically grounded phenomenon. Through concrete examples—such as temperature, magnetism, and herd immunity in social networks—we review how collective behavior can arise from local interactions that are constrained by global boundaries. By refining the concept of emergence, we gain a clearer and more grounded understanding of complex systems. Our goal is to show that emergence, when properly framed, offers not mysticism but insight.
@@
This manuscript was compiled on October 7, 2025
@@
Emergence occurs where a whole exhibits properties absent from its individual parts—an idea that traces back to Aristotle’s dictum that “the whole is something besides the parts” (1). In science, the term refers to patterns, structures, or behaviors that are not straightforwardly deducible from components in isolation (2, 3, 4, 5, 6, 7), yet are often captured by coarse, predictive descriptions that do not track every micro-detail. You can see emergence in the wild—flocks wheeling as one, fireflies locking into rhythm, insect societies coordinating without commanders (8)—and through computer simulations, where minimal rules produce striking order: Reynolds’ “boids” (9), cellular automata (10), and the Game-of-Life glider (11).
Philosophical roots of the concept trace back to Mill’s distinction between additive “mechanical” effects and qualitatively new “chemical” ones (12). Lewes later coined the term “emergent” for such cases (13), and early British emergentists (14, 15, 16, 17) developed it into a layered view of nature in which higher-level laws complement, and sometimes influence, lower-level processes. This tradition faded as reductionism gained ground, driven by the intertheoretic reduction models (18, 19, 20). The debate revived with arguments for multiple realizability (21, 22, 23) and Anderson’s claim that “more is different” (24), followed by work on supervenience and causal autonomy (25). In recent years, the term has evolved from describing specific transitions, such as the “emergence of metabolism” (26), to becoming a popular buzzword, gaining more usage in scientific and non-scientific discourses (27, 24). Fig. 1 shows this trend.
Here, we review operational perspectives on emergence—consistent with physicalism and grounded in effective theories and information-theoretic principles—rather than non-mainstream viewpoints (28, 29, 30, 31, 32, 33, 34). In the sections that follow, we assemble the concepts required for a working account of emergence across physics, biology, and the social sciences. We formalize emergence through lossy compression that preserves the information relevant for prediction and control, and we motivate Effective Theories as the mechanism that secures autonomy and multiple realizability. We then treat quantitatively secure testbeds—critical phenomena and phase transitions—to show how order parameters, response functions, scaling, and data collapse locate the onset of emergence and render it mechanistic and testable. Next, we use dualities to show that distinct microdescriptions can yield the same macrophysics, grounding autonomy without extra causes. Finally, we distinguish between explanatory and ontological reduction: reduction secures cross-scale consistency, while effective theories deliver explanations and control. Sequenced this way, the paper justifies an economy-of-description perspective: we can explain how many emergent phenomena happen—even if not yet all—and none require supernatural causes.
@@
Levels of Description
Knowledge unfolds across successive levels of abstraction—from raw experience to language and higher-order concepts (36). Confusion between these levels distorts understanding. Emergence is everywhere, and the surprise it evokes often reflects where we choose to stop asking more questions (37, 25). What we call “emergent” typically depends on the limits of our knowledge, tools, or perspective—in other words, on epistemology (38, 39). Mixing red and green paint yields a dull brown—basic chemistry and optics explain it well enough for most, and few would call the result emergent. Yet a physicist who tracks the same blend down to quantum‐mechanical interactions would hardly find the task trivial.
Explanation is always a matter of perspective: an economist analyzing a financial crisis need not invoke atoms or molecules but instead works at the social scale relevant to the questions at hand. Every phenomenon can be analyzed at multiple descriptive levels, and what feels emergent at one level may be obvious at another (38). Even so, emergent behaviors are not merely in the eye of the beholder—they can display objective, quantifiable signatures such as information flow or causal strength (40, 41). We can make emergence more rigorous by formalizing it through coarse-graining maps that discard detail yet preserve prediction.
@@
Emergence is present when there exists a many-to-one map from a micro-level theory (more fundamental, detailed, or lower-level) to a macro-level theory (higher-level), such that the macro description remains predictive even after discarding most of the microscopic detail (42). As Fig. 2(a,b) presents, a map is a selective, structured, purpose-driven abstraction—useful precisely because it is not the whole territory it depicts (36, 43). Every map omits, distorts, or abstracts; acknowledging those sacrifices is part of scientific hygiene (43).
Coarse-graining is the process of identifying which features of a system are essential for capturing its macroscopic behavior and which can be systematically discarded. Therefore, it leaves us with just enough structure and real patterns to construct an autonomous and useful theory (44). This is why engineers who build bridges do not need to take a course in quantum field theory, or why airplanes can fly safely despite our incomplete understanding of quantum gravity. Suppose a system’s state at time $t$ is given by the vector $\mathbf{x}(t)=\bigl(x_{1}(t),\dots,x_{N}(t)\bigr)$, where a high-resolution dataset comprises $N$ time series, and each $x_{i}(t)$ represents the evolution of the $i$-th microscopic variable. The coarse-graining can be described as a mapping:
@@
where $N^{\prime}<N$, and the coarse-grained observables $\{X_{j}\}$ may be defined on a different (typically coarser) time grid $\{t^{\prime}_{\ell}\}$. Therefore, $\mathcal{F}$ is a lossy data compression where only the information needed for prediction survives across scales (45, 29). Fig. 2 pictures such maps. This viewpoint offers a substrate-independent vocabulary for linking small-scale descriptions to large-scale laws. In high-energy physics, the information relevant for prediction determines which effective theory you use at each scale (46). In brains and societies, the same approach helps identify mesoscales that are maximally compressed yet still predictive (47, 48, 49, 50, 51). Panels (a) and (b) show two different sets of coarse-graining where they average out the observables. Panel (e) illustrates how, by compressing data through the right process, we arrive at a simpler theory. The map $\mathcal{F}$, however, need not be simple, trivial, local, or unique.
The first step toward understanding how emergence occurs is to identify when and under what conditions it arises, as well as its immediate consequences. Useful large-scale descriptions may need entirely different degrees of freedom than the ones appearing in the small-scale theory, and they can realize symmetries in new ways. Choosing the right variables is a scientific discovery step, not a fixed recipe. The emergence of life, mind, or social structure appears puzzling not because they conflict with physicalism, but because they represent a qualitatively more complex class of emergent phenomena (52). Therefore, there is no universal recipe for identifying the “right” macroscopic variables. The key practical questions here are whether there exists a finite set of macroscopic parameters that can adequately describe the phenomena of interest within a specified tolerance, and if so, which mapping $\mathcal{F}^{\star}$ minimizes the chosen distortion measure while achieving that macro-level description. The outcome is a systematic and metric-aware approach to transition from micro-level dynamics to macro-level dynamics.
Figure 2: Coarse-graining, Universality, and RG-flow. (a) Temperature as a coarse-grained variable: under local equilibrium, averaging kinetic energies defines temperature, $\mathcal{F}:(v_{1},\ldots,v_{N})\!\mapsto\!T$. (b) Local coarse-graining by $2{\times}2$ blocks with majority-vote, $S_{I}=\operatorname{sgn}(s_{i}+s_{j}+s_{k}+s_{l})$ for $s_{\ell}\in\{\pm 1\}$; iterating the map flows toward a uniform block-spin configuration. (c) Liquid–gas criticality shares the Ising universality class. The control parameter is the reduced temperature $\tau=(T-T_{c})/T_{c}$. On a lattice, the lattice-gas mapping $n_{i}\!\in\!\{0,1\}\mapsto\sigma_{i}=2n_{i}-1$ (occupied $\uparrow$, empty $\downarrow$) makes the density deviation $\Delta\rho=\rho-\rho_{c}$ directly analogous to the magnetization $m$. (d) Increasing temperature drives a continuous transition from an ordered phase to a disordered phase, and the transition point we see scaling behaviors (28, 53). (e) Coarse-graining compresses high-resolution data and induces a renormalization-group (RG) flow in theory space. When the coarse-grained description is related to the microscopic one by parameter redefinitions, the model is said to be renormalizable. A macro-law screens off micro-details but remains autonomous, meaning that macro variables aren’t just convenient summaries but “real enough” to be the subject of genuine laws and explanations—even if many different microstates realize the same macrostate.
@@
The very notion of a map between levels of description tempts us to believe that deeper, more fundamental layers offer better explanations—an intuition reinforced by Physics’ success in reducing systems to atoms and beyond. But how far should we go? What counts as the most fundamental level? Physics is expressed in intrinsically redundant languages—coordinates, gauge potentials, duality frames, and so on—and those redundancies shift as we move between scales (54). What we call “fundamental”—even spacetime and gauge fields—may not exist in a more microscopic theory, but rather emerge after coarse-graining. Spacetime geometry and general relativity might also be emergent, arising from quantum theory through entanglement and boundary dynamics (54). This realization, however, does not make emergence a “secret sauce.” It simply shows how nature organizes complexity through scale, and “fundamental” is often a provisional concept.
@@ A New Ontology
Classical Mechanics itself may emerge from a quantum theory, which is, in turn, ontologically different (61). In Classical Mechanics, a particle is represented by a point in phase space with well-defined position and momentum. In Quantum Mechanics, the ontology shifts: a particle is represented by a vector in Hilbert space, and is described by a wavefunction (62). Position and momentum are no longer coordinates but observables. The uncertainty principle ensures that no wavefunction can define both position and momentum with perfect precision. Thus, quantum mechanics offers no definitive answer to the classical question of “where” a particle is, or “how fast” it is moving (62). Nevertheless, the classic macroscopic theory remains applicable in its domain, and it is consistent with the underlying quantum description through the processes of decoherence and measurement, which effectively produce the appearance of wavefunction collapse (63).
@@
Since the advent of Statistical Mechanics, we have learned how coarse-graining microscopic degrees of freedom gives rise to macroscopic quantities, such as temperature or pressure, and how these quantities are connected through equations of state (64). Statistical mechanics not only explains why thermodynamics works but also defines its domain of applicability. For example, the domain of applicability of temperature can be expressed as
@@
where $m$ is the particle mass, $\langle v^{2}\rangle$ is the ensemble-averaged squared speed, $N$ is the number of particles, and the Boltzmann constant is set to one, $k_{B}=1$ (28). For large $N$, the fluctuation term $1/\sqrt{N}$ vanishes, making temperature a sharply defined coarse-grained variable. For small systems, however, fluctuations are significant, and temperature loses descriptive power.
@@
In the large-size limit, Thermodynamics is not merely an approximation to Statistical Mechanics—it is an effective theory (65), one that captures the essential behavior of macroscopic systems using a few key variables at the right scale (66, 67). Effective theories isolate what matters at a given scale and discard what does not (45). They are not shortcuts; they are self-contained, predictive, and often universal descriptions of emergent levels of reality. That is why, for specific questions, instead of tracking the detailed positions and velocities of every particle, we use variables like temperature, pressure, and entropy with their own simple and powerful macro-level laws (68, 69). A mean-field theory is the zeroth-order effective description (70). It is invaluable as a first cut, exact when fluctuations vanish, but insufficient whenever correlations hold the key to collective behavior (71).
@@
Not all effective theories are created equal. In some systems—particularly those involving biological, cognitive, or social dynamics—the coarse-grained variables do not merely summarize the current state of the underlying components; they also retain memory, storing information about past configurations. Such memory and nonlocality are not exclusive to living or social systems: in physics, integrating out fast fields generically produces spatially nonlocal interactions and temporal memory kernels (72, 73, 74). Even gravity has been cast in explicitly nonlocal, memory-like form (75, 76, 77). This historical dependence alters the structure of the effective theory, introducing higher-order terms that reflect feedback, path dependence, and self-reference (42). Such systems are still governed by physical laws, but their dynamics can become computationally intractable, meaning that no tractable microscopic derivation will fully recover the macroscopic behavior (52).
Onset of Emergence, Symmetry Breaking & Criticality
@@
A clear and intuitive example of how emergence happens is magnetization. In a fridge magnet, the collective alignment of billions of electron spins (little magnets) produces a macroscopic magnetic property, denoted $m(T)$, which appears only when the system is below a critical temperature $T_{c}$. For a fridge magnet, it is around $450^{\circ}$C. Heating the material beyond this point destroys the alignment, and as Fig. 2d shows, the magnetic property disappears in a continuous phase transition (28). In other words, heating injects fluctuations that destroy spin alignment, driving the system from a magnetically ordered phase ($m>0$) to a disordered paramagnetic phase ($m=0$). A single governing equation (Ising Hamiltonian) describes both magnetized and non-magnetized states (28).
At the critical point $T=T_{c}$, the spin-flip symmetry of the Hamiltonian remains intact, but the actual state chooses a specific direction, resulting in $m\neq 0$ (70, 78). This symmetry breaking defines an order parameter that compresses the full $10^{23}$-spin microstate into a single coarse-grained vector whose dynamics—like domain walls or spin waves—obey new effective laws (79, 80, 24). At $T_{c}$, key properties of the system—such as correlation length and relaxation time—diverge (or peak in finite systems), with finite-size scaling and data collapse as empirical checks (81, 82, 28). These divergences, characterized by critical exponents, serve as clear markers of the onset of emergence.
More interestingly, different systems composed of distinct elements can exhibit the same critical behavior (83). That is, systems with very different microstructures can fall into the same universality class and be described by the same macroscopic theory. This substrate-independence also means that, from the perspective of the emergent behavior, one cannot tell which microscopic system produced it. When very different microscopic models display the same large-scale behavior, even exceptionally precise measurements of large-scale behavior usually cannot identify a unique microscopic theory without extra assumptions (e.g., symmetry content or locality) (84, 81).
@@
The liquid–gas transition belongs to the same universality class as the ferromagnetic transition (53). Fig. 2c illustrates the technique of translating these two transitions into one another. Universality classes, therefore, organize not just critical exponents but also the very symmetry content that survives at macroscopic scales, reinforcing why broken symmetry is one maker of emergence (24, 54). In this sense, “more is different” (24): once many degrees of freedom lock together, they generate effective laws—and sometimes entirely new effective symmetries—that are absent from, yet fully compatible with, the underlying dynamics.
It is worth mentioning that information-theoretic diagnostics similarly pinpoint the onset of emergence: mutual information rises near criticality, transfer entropy reveals increasingly predictive interactions, and persistent mutual information isolates long-lived choices beyond short-time correlations (45, 49, 50). Causal tests complement these signals: Effective Information can increase under coarse-graining—so a well-chosen macro description exerts greater causal constraint than the micro—and partial information decomposition exposes downward causation and causal decoupling directly in multivariate data (85, 86).
Renormalization & Universality
@@
Technically, critical universality manifests only within the critical region of a continuous phase transition: as control parameters approach the transition, the renormalization-group (RG) flow is drawn toward a scale-invariant fixed point—attractive along irrelevant directions and unstable along relevant ones—so macroscopic behavior becomes largely independent of microscopic details (84, 87, 88). This picture applies to both equilibrium and nonequilibrium continuous transitions (89, 90); nothing forbids far-from-equilibrium universality (91), though for certain many-body Hamiltonians even basic features (spectral gap, phase diagram, etc.) are undecidable in principle despite a computable RG flow (92).
A canonical non-equilibrium exemplar is the class of absorbing-state transitions (93, 94, 95): dynamics are irreversible, the system can enter configurations it cannot escape, and the order parameter (e.g., the active-site density in the contact process) vanishes continuously at a critical control parameter with power-law correlations at criticality (89, 90). Under broad conditions, these phase transitions fall into the directed percolation (DP) universality class, often regarded as the non-equilibrium analogue of the Ising paradigm we mentioned before (89, 90). Neural-population recordings, deep neural networks, and insect swarms also exhibit RG-style scaling in nonequilibrium settings (96, 97, 98, 99, 100, 101). What truly matters for RG are locality, symmetry, and conservation constraints, and dimensionality (89, 102). That said, beyond a few well-understood universality classes, we still lack a clear, well-founded theoretical framework—comparable in scope to the renormalization group—for systematically organizing and classifying non-equilibrium phases (89).
One must take care to remember that criticality is neither the only route to emergence nor trivial to establish empirically (103). For example, it is not straightforward to extend the statistical mechanics framework of critical phenomena to describe abrupt qualitative changes in dynamical systems. Likewise, the emergence of life from chemical interactions does not neatly correspond to a phase transition in the physical sense, despite claims to the contrary by some physicists (104). Intricate phenomena such as consciousness or thought could, in principle, arise from entirely different physical substrates—biological neurons or silicon circuits alike (105, 106)—provided there is sufficient evidence that they result from well-defined critical phenomena (107). Across biological systems, current evidence favors near-criticality over exact self-tuned criticality (108, 109, 110). In neuroscience in particular, recent surveys argue that a distance-to-criticality perspective is more realistic and useful—for cognition and pathology—than the claim that the brain sits precisely at criticality; both edge-of-chaos and avalanche-like regimes can appear, often state-dependently (108, 109, 110). Therefore, “critical-looking” statistics (power laws, avalanches) are insufficient on their own (111, 96, 112, 113, 114) and broken symmetries underpin not all emergent behaviors (115, 116). Accordingly, strong claims of criticality, let alone emergence in living systems, require careful analysis (82, 117, 28, 89).
Duality & Emergence
Duality is an equivalence (often infrared-exact (118, 119)) between two seemingly different theories: a one-to-one “dictionary” that maps observables, states, and parameters so that all physical predictions coincide. Dualities reveal that theories with ontologically distinct variables—such as bosons versus fermions, or gravity versus no gravity—can describe the same underlying physics (118). Formally, a duality establishes a structural correspondence between the measurable quantities of two theories (a *-isomorphism) such that every observable, expectation value, and response function matches once external conditions—sometimes called “sources”—are translated accordingly. This agreement may hold up to local adjustments, like contact terms or counterterms, reflecting different renormalization schemes (118).
In 1+11{+}1 dimensions, the bosonic sine–Gordon model is exactly dual to the fermionic Thirring model, exchanging solitonic waves for interacting particles and relating their couplings nonlinearly (120). In the Anti–de Sitter/conformal field theory correspondence, a conformal field theory without gravity on a $d$-dimensional boundary is holographically equivalent to a (d+1)(d{+}1)-dimensional bulk theory with gravity (46). Classic examples from statistical and gauge theories—such as high–low temperature mappings, electric–magnetic duality, and particle–vortex correspondence—further illustrate how distinct microscopic descriptions can encode the same macroscopic behavior (121, 122, 123, 124)
What we learn from dualities is twofold: First, large-scale behavior is not dictated by microscopic “stuff” (but by invariants preserved under the duality map—correlators, partition functions, anomalies, and consistency with locality and boundary conditions) (119). Second, because dualities often relate weak and strong coupling, they reveal multiple, equally valid languages for describing the same phenomena (119). Macroscopic structure, therefore, can appear as alternative readouts of a single physical substrate. Which language proves most useful depends on the question at hand—computational tractability, available probes, or symmetry constraints—rather than on a uniquely privileged ontology.
@@
Emergence in Networks & Social Systems
Moving beyond physics, we consider how large-scale patterns emerge from local interactions in social and informational systems. Emergent phenomena possess causal power. In both physical and social worlds, entities like society and culture act as real forces shaping outcomes. Even in systems designed without formal rules—such as decentralized communities like Wikipedia or Stack Exchange—structures, hierarchies, and power distributions inevitably emerge, and the resulting social laws profoundly influence the world.
@@
Consider a set of $N$ points where each pair is connected with probability $p$. This defines an Erdős–Rényi random graph with $N$ nodes and roughly $p\binom{N}{2}$ links or connections. As $p$ increases, the network shifts from isolated nodes ($p=0$) to a fully connected graph ($p=1$), with clusters forming in between. Given the value of $p$, it is interesting to look at the behavior of the cluster with the most number of connected nodes, the giant component. Letting $\sigma$ be the fraction of nodes it contains, we find that as in the large-size limit ($N\to\infty$), the network undergoes a phase transition such that the size of its giant component is zero below a critical value $p_{c}$, and positive above it. Here, $p$ is the control parameter, and $\sigma$ represents the order parameter, mirroring magnetization in ferromagnets, which appears below the critical temperature. When $p>p_{c}$, large-scale connectivity emerges, and the network is said to percolate (125).